Anna's mother will in 12yrs be three times as old as she was 20yrs ago, how old is Anna's mother now
Answer by ankor@dixie-net.com(22740)

Anna's mother will in 12yrs be three times as old as she was 20yrs ago, how old is Anna's mother now
:
m + 12 = 3(m-20)
m + 12 = 3m - 60
12 + 60 = 3m - m
72 = 2m
m = 36 yr now
:
:
Check this
36 + 12 = 3(36-20)
48 = 3(16)